# Superposition Yields Robust Neural Scaling

### v2

@@ 5 Discussion
We acknowledge that our work is built on observations of the toy model and shallow theoretical analysis without rigorously solving the toy model. We are thus limited to explaining some deeper questions. For example, Equation (7) is an observation. Although we explained the two extremes, $\alpha_{m}=1$ and $\alpha_{m}=2(\alpha-1)$, and provided heuristics for a transition, we cannot predict the exact transition.
Neural scaling laws also include scaling laws with dataset size and with training steps, which we did not study. At each step, a fixed number of new data are used for optimization. So, we expect the scaling with the total data amount and that with training steps will be the same, similar to the results at weak superposition [19]. However, in the strong superposition regime, data or training step scaling is related to angle distribution and how angles between representations evolve, which cannot be easily explained without rigorous solving. Future works with better theoretical analysis of our toy model are needed to answer these detailed or extended questions.
Our analysis suggests LLMs are in the strong superposition regime, but the underlying reasons were not studied. Specifically, LLMs do not need weight growth $\gamma<0$ to have strong superposition. The use of cross-entropy loss or softmax function may be important — minimizing the interference from an unimportant token prefers to make the representation non-zero with negative overlaps with others. Another contributing factor may be that $m^{2}/2>n$ in LLMs, which could make it easier for a nearly mutually orthogonal configuration to form, thereby favoring superposition.
We focused on representation loss in this study, yet LLMs should also have losses due to parsing or processing in the transformer layers. As a conceptual discussion, we imagine that the loss associated with model size can be written as
@@
where $\ell$ is the depth of the LLM, $f_{m}$ and $f_{\ell}$ are two functions capturing the loss due to representation and parsing, respectively. We can write the equality because $\ell$ depends on $m$ in LLMs [38, 39, 40, 41]. Given model size $N$, $m$ and $\ell$ are constrained (roughly, $N\propto m^{2}\ell$). There is an optimal $m$$-\ell$ relationship such that the loss $f_{m}(m)+f_{\ell}(\ell)$ can be minimized given $N$ [54]. At this optimal $m$$-\ell$ relationship, $f_{m}(m)$ and $f_{\ell}(\ell)$ should be balanced. One can see that the empirical $m$$-\ell$ relationship is not optimal according to [54] but is close. Therefore, following such $m$$-\ell$ relation, we expect $f_{\ell}(\ell)$ to be similar to $f_{m}(m)$. And if $f_{m}(m)\sim 1/m$ due to superposition and $f_{\ell}(\ell)$ is similar, we can measure an empirical $\alpha_{m}\approx 1$ from data, which is true. So, we conclude that superposition is an important mechanism underlying the loss and scaling laws we see. However, since we do not know the function $f_{\ell}(\ell)$, we cannot predict the optimal $m$$-\ell$ relation or the exponent $\alpha_{N}$.
A natural future direction is to study the parsing-limited scaling (i.e, $f_{\ell}(\ell)$ function). We hypothesize that parsing-limited scaling may resemble the scaling behavior observed when fitting functions or manifolds. The open-source LLMs we examined, as well as those studied in prior neural scaling law papers [2, 3], span model sizes from 100M to 70B parameters, where width (model dimension) varies. However, in the regime from 70B to 700B [38, 39, 40, 41], model scaling seems to focus on increasing depth. We therefore expect parsing-limited scaling to become important for models larger than 70B. It is also plausible that the observed scaling of inference time [55] is connected to this parsing-limited regime. The study of neural scaling laws can be refined by distinguishing between width-limited and depth-limited regimes. In each regime, there should be loss decay behaviors with model size, dataset size, and training steps, highlighting the need for further investigation.
@@
Beyond explaining existing phenomena, our results may offer guidance for future LLM development and training strategies. If our framework accounts for the observed neural scaling laws [3], we suggest that this kind of scaling is reaching its limits — not because increasing model dimension is impossible, but because it is inefficient. Recognizing that superposition benefits LLMs, we propose that encouraging superposition could enable smaller models to match the performance of larger ones and make training more efficient. Recent architectures such as nGPT [56], which constrain hidden states and weight matrix rows to lie on the unit sphere, demonstrate improved performance and faster training. We interpret these methods as analogous to our weight growth mechanism, both of which promote superposition. Conversely, weight decay is expected to suppress superposition. Prior work has shown that weight decay in LLMs is used primarily for training stability rather than performance gains. Alternative optimization techniques that stabilize training without weight decay have shown promising results [57], potentially due to enhanced superposition. In summary, promoting strong superposition may be a fruitful direction for the design of both LLM architectures and optimization algorithms.
@@
As a side note, while pre-training loss is a key indicator of model performance, it should not be the sole metric of interest. At the same loss level, LLMs with different degrees of superposition may exhibit differences in emergent abilities such as reasoning or trainability via reinforcement learning [58]. Understanding these differences is an important direction for future research.
@@
In conclusion, we demonstrated that superposition of representations can give rise to fast power law decay of loss with model size — a phenomenon that holds across a wide range of data structures and may help explain the neural scaling laws observed in LLMs [3]. Our results contribute to a deeper understanding of modern artificial intelligence systems. This work also opens rich directions for future research. We hope our insights will support the continued development and training of more capable and efficient LLMs.
@@ A.1 Toy model loss
We now review our proposed configuration, consisting of strongly represented and weakly represented features. The first $\phi_{1}n$ most important features are considered to be strongly represented, whose overlap with any other representation scales as $\sqrt{1/m}$. The rest of the features are weakly represented and squeezed into a small angle such that they have small overlaps with the strongly represented, while they can have large overlaps with each other. With such a configuration, the first $\phi_{1}n$ terms in the summation of Equation (10) will scale as $1/m$ since each term $\langle\cdots\rangle_{v_{i}}$ scales as $1/m$. The rest of the terms in Equation (10), in the worst scenario that the terms $\langle\cdots\rangle_{v_{i}}$ do not decrease obviously with $m$, will be proportional to $\sum_{i=\phi_{i}n}^{n}p_{i}$. If the strongly represented features dominate, we can have $1/m$ scaling for the loss. On the contrary, when the weakly represented dominate, we expect the loss to have scaling like $\sum_{i=\phi_{i}n}^{n}p_{i}$. More specifically, if $p_{i}\sim 1/i^{\alpha}$ ($\alpha>1$) and we use $m^{2}/2$ to approximate $\phi_{i}n$, the loss scales as $1/m^{2(\alpha-1)}$.

### v3

@@ 5 Discussion
We acknowledge that our work is built on observations of the toy model and shallow theoretical analysis without rigorously solving the toy model. We are thus limited to explaining some deeper questions. Our analysis suggests LLMs are in the strong superposition regime, but the underlying reasons were not studied. The use of cross-entropy loss or the softmax function may be important. And features are quite sparse in language, as the number of words in each sentence is much fewer than the vocabulary size, thereby favoring superposition.
Neural scaling laws also include scaling laws with dataset size and with training steps, which we did not study. At each step, a fixed number of new data points are used for optimization. So, we expect the scaling with the total data amount and that with training steps will be the same, similar to the results at weak superposition [20]. However, in the strong superposition regime, data or training step scaling is related to angle distribution and how angles between representations evolve, which cannot be easily explained without rigorous solving.
We focused on representation loss, yet LLMs should also have losses due to parsing or processing in the transformer layers. We imagine that the loss associated with model size can be written as
@@
where $\ell$ is the depth of the LLM, $f_{m}$ and $f_{\ell}$ are two functions capturing the loss due to representation and parsing, respectively. A future direction is to study the parsing-limited scaling (i.e, $f_{\ell}(\ell)$ function) independently. It is also plausible that the observed scaling of inference time [58] is connected to this parsing-limited regime. We here write the equality because $\ell$ depends on $m$ in LLMs [39, 40, 41, 42]. Given model size $N$, $m$ and $\ell$ are constrained (roughly, $N\propto m^{2}\ell$). There is an optimal $m$-$\ell$ relationship such that the loss $f_{m}(m)+f_{\ell}(\ell)$ can be minimized given $N$ [59]. At this optimal $m$-$\ell$ relationship, $f_{m}(m)$ and $f_{\ell}(\ell)$ should be balanced. Therefore, we expect $f_{\ell}(\ell)$ to be similar to $f_{m}(m)$. And if $f_{m}(m)\sim 1/m$ due to superposition and $f_{\ell}(\ell)$ is similar, we can measure an empirical $\alpha_{m}\approx 1$ from data, which is true. Or, if the width-limited loss is much larger, we can also observe that the total loss due to model size has $\alpha_{m}\approx 1$. We conclude that superposition in any case is an important mechanism underlying neural scaling laws.
Beyond explaining existing phenomena, our results may offer guidance for future LLM development and training strategies:
Assuming our explanation of width scaling is correct, we ask can we change the loss scaling with width to be faster than power laws, or to have larger exponents? The answer is no for natural languages but may be yes for domain tasks with super skewed feature frequencies. Another question is when the scaling law will stop? Based on our naive connection between features and tokens, the answer is that when the model dimension reaches the vocabulary size, the loss limited by width will deviate from a power law and vanish. However, the vocabulary size may set a lower bound for the true number of independent things in language, then the power law with width may continue for a longer time.
@@
Recognizing that superposition benefits LLMs, encouraging superposition could enable smaller models to match the performance of larger ones (with less superposition) and make training more efficient. Architectures such as nGPT [60], which constrain hidden states and weight matrix rows to the unit sphere (promoting superposition), demonstrate improved performance. Optimizers that stabilize training without weight decay have also shown promising results [61], potentially due to enhanced superposition. Yet, these improvements may be related to altering coefficients in the neural scaling laws rather than the exponents. We also acknowledge that encouraging superposition may cause difficulties for the mechanistic interpretation of models and AI safety [27, 62].
@@
As a side note, with the same pre-training loss, LLMs with different degrees of superposition may exhibit differences in emergent abilities such as reasoning or trainability via reinforcement learning [63], requiring future studies.
@@
In conclusion, we studied when loss can be a power law and what the exponent should be with different data properties and degrees of superposition. We found that geometric interference at strong superposition may explain the LLM neural scaling laws observed [3]. Our results contribute to a deeper understanding of modern artificial intelligence systems, which also open various directions for future research. We hope our insights will support the continued development and training of more capable and efficient LLMs.
Acknowledgments and Disclosure of Funding
We are grateful for feedback and suggestions from Yasaman Bahri, Cengiz Pehlevan, Surya Ganguli, Blake Bordelon, Daniel Kunin, and the anonymous reviewers. The authors acknowledge the MIT Office of Research Computing and Data for providing high performance computing resources that have contributed to the research results reported within this paper. J. G. thanks the Sloan Foundation for funding. The authors declare no competing interests.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims match theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We start our Discussion section by discussing limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Assumptions are clearly stated.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all details sufficient to reproduce.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We submit our code in the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We explained these details in our Appendices.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We included error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide this information in the Appendices.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We followed the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work is theoretical, with no obvious social impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not have such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cited the datasets and LLMs analyzed.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ A.1 Toy model loss
For convenience, based on the observation that the vector norms are bimodal around $1$ in the strong superposition regime, we define strongly represented features as those that have $\|W_{i}\|_{2}>1$, which are more frequent and ETF-like, and weakly represented ones for those with $\|W_{i}\|_{2}<1$. We can quantify the fraction of strongly represented as
$\phi_{1}=|\{i:\|W_{i}\|_{2}>1\}|/n,$ (9)
which significantly exceeds $m/n$ and is around the ETF expectation $m^{2}/2n$ (Figure 16).
We now review our conjectured extreme configuration, consisting of strongly represented and weakly represented features. The first $\phi_{1}n$ most important features are considered to be strongly represented, whose absolute overlap with any other representation scales as $\sqrt{1/m}$. The rest of the features are weakly represented and squeezed into a small angle such that they have small overlaps with the strongly represented, while they can have large overlaps with each other. With such a configuration, the first $\phi_{1}n$ terms in the summation of Equation (8) will scale as $1/m$ since each term $\langle\cdots\rangle_{v_{i}}$ scales as $1/m$. The rest of the terms in Equation (8), in the worst scenario that the terms $\langle\cdots\rangle_{v_{i}}$ do not decrease obviously with $m$, will be proportional to $\sum_{i=\phi_{i}n}^{n}p_{i}$. If the strongly represented features dominate, we can have $1/m$ scaling for the loss. On the contrary, when the weakly represented dominate, we expect the loss to have scaling like $\sum_{i=\phi_{i}n}^{n}p_{i}$. More specifically, if $p_{i}\sim 1/i^{\alpha}$ ($\alpha>1$) and we use $m^{2}/2$ to approximate $\phi_{i}n$, the loss scales as $1/m^{2(\alpha-1)}$.